OpenStudy (bethgrace96):

evaluate the function x=0 and x=1 and x=2. then write a recursive formula for the pattern. f(x)=250(0.5)^x. WILL GIVE MEDAL FOR FIRST COMPLETE ANSWER

OpenStudy (campbell_st):

well let x = 0 so \[f(0) = 250 \times 0.5^0~~~~so~~~f(0) = 250\] \[f(1) = 250 \times 0.5^1~~~so ~~f(1) = 125\] lastly \[f(2) = 250 \times 0.5^2 = 62,5\] so you have a geometirc series 250, 125, 62.5 so the common ratio is 0.5 and the 1st term is 250... hope it helps you get the recursive equation